# Knots and variance ordering of sequential Monte Carlo algorithms

### v1

@@ 6.1 Extended models
6.2 Terminal knots
Definition 6.3 (Terminal knot compatibility).
  1. (i)
    The model satisfies $M_{n}^{\circ}=U\otimes V^{G_{n}\cdot\phi}$ and $G_{n}^{\circ}=V(G_{n}\cdot\phi)\otimes\phi^{-1}$, for some Markov kernels $U$ and $V$, reference model $\mathcal{M}=(M_{0:n},G_{0:n})$, and target function $\phi$.
  2. (ii)
    The knot satisfies $U=RK$.
Overall, Definition 6.3 extends the notion of knot-model compatibility to the special case of terminal knots, and hence expands the compatible knot-model set, $\mathscr{D}_{n}$, given in (5). Recall that the knot operator, $\ast:\mathscr{D}_{n}\rightarrow\mathscr{M}_{n}$, maps compatible knot-model pairs to the space of Feynman–Kac models for horizon $n\in\mathbb{N}_{0}$. Definition 6.4 extends this operation to terminal knots.
Definition 6.4 (Knot operator, terminal knots).
Consider a terminal knot $\mathcal{K}=(n,R,K)$ and model $\mathcal{M}=(M_{0:n},G_{0:n})$. If $M_{n}=P_{1}\otimes P_{2}$ and $G_{n}=H\otimes\phi^{-1}$ then the knot operation yields $\mathcal{K}\ast\mathcal{M}=(M_{0:n}^{\ast},G_{0:n}^{\ast})$ where
$M_{n}^{\ast}=R\otimes K^{H}P_{2},\quad G_{n}^{\ast}=K(H)\otimes\phi^{-1},$
Note that the existence of $P_{1}$ $P_{2}$, $H$, and $\phi$ in Definition 6.4 are guaranteed by compatibility condition (i). However, this still leaves the question of what models have the form to satisfy the model compatibility condition. Taking $U=M_{n}$ and $V=\mathrm{Id}$, demonstrates that a $\phi$-extended model has the correct form, and the general case is presented next.
Proposition 6.5 (Form of extended models with knots).
Consider a model $\mathcal{M}^{\circ}\in\mathscr{M}_{n}$. For some $m\in\mathbb{N}_{1}$, if there exists a sequence of knots $\mathcal{K}_{1},\mathcal{K}_{2},\ldots,\mathcal{K}_{m}$ such that $\mathcal{M}^{\circ}=\mathcal{K}_{m}\ast\cdots\ast\mathcal{K}_{2}\ast\mathcal{K}_{1}\ast\mathcal{M}^{\phi}$ for some reference model $\mathcal{M}=(M_{0:n},G_{0:n})$ and target function $\phi$ then there exists Markov kernels $U,V$ such that $M_{n}^{\circ}=U\otimes V^{G_{n}\cdot\phi}$ and $G_{n}^{\circ}=V(G_{n}\cdot\phi)\otimes\phi^{-1}$.
@@ 6.2 Terminal knots
By assumption $\mathcal{M}^{\circ}=\mathcal{K}_{m}\ast\cdots\ast\mathcal{K}_{2}\ast\mathcal{K}_{1}\ast\mathcal{M}^{\phi}$ for knots $\mathcal{K}_{1},\mathcal{K}_{2},\ldots,\mathcal{K}_{m}$. First consider the $\phi$-extended model $\mathcal{M}^{\phi}=(M_{0:n}^{\phi},G_{0:n}^{\phi})$. Take $U=M_{n}$ and $V=\mathrm{Id}$ and noting Definition 6.1 we can conclude that $\mathcal{M}^{\phi}$ satisfies the required form.
Let $\mathcal{M}_{s+1}=\mathcal{K}_{s}\ast\mathcal{M}_{s}$ for $s\in[1\,{:}\,m]$ where $\mathcal{M}_{1}=\mathcal{M}^{\phi}$, and note that $\mathcal{M}_{m+1}=\mathcal{M}^{\circ}$}. For some $t\in[1\,{:}\,m]$ suppose $\mathcal{M}_{t}=(M_{t,0:n},G_{t,0:n})$ is $\phi$-accordant with $\mathcal{M}$ then we have $M_{t,n}=U_{t}\otimes V_{t}^{G_{n}\cdot\phi}$ and $G_{t,n}=V_{t}(G_{n}\cdot\phi)\otimes\phi^{-1}$ for some Markov kernels $U_{t}$ and $V_{t}$. Now consider $\mathcal{M}_{t+1}=\mathcal{K}_{t}\ast\mathcal{M}_{t}=(M_{t+1,0:n},G_{t+1,0:n})$.
First case. If $\mathcal{K}_{t}$ is a non-terminal knot or a knotset then $M_{t+1,n}=K^{G_{n-1}}U_{t}\otimes V_{t}^{G_{n}\cdot\phi}$ and $G_{t+1,n}=V_{t}(G_{n}\cdot\phi)\otimes\phi^{-1}$ for some Markov kernel $K$. Letting $U_{t+1}=K^{G_{n-1}}U_{t}$ and $V_{t+1}=V_{t}$ shows that $\mathcal{M}_{t+1}$ satisfies the required form.
Second case. If $\mathcal{K}_{t}$ is a terminal knot then $M_{t+1,n}=R\otimes K^{V_{t}(G_{n}\cdot\phi)}V_{t}^{G_{n}\cdot\phi}$ and $G_{t+1,n}=KV_{t}(G_{n}\cdot\phi)\otimes\phi^{-1}$ for some Markov kernels $R$ and $K$ such that $RK=U_{t}$. By Proposition A.3, we can state $M_{t+1,n}=R\otimes(KV_{t})^{G_{n}\cdot\phi}$ and hence letting $U_{t+1}=R$ and $V_{t+1}=KV_{t}$ shows that $\mathcal{M}_{t+1}$ satisfies the required form.
Therefore, by induction, any model can be represented as a $\phi$-extended model with knots has the required form. ∎
Typically, the first application of a terminal knot will be on a $\phi$-extended model, which we demonstrate in Example 6.6.
Example 6.6 (Terminal knot for $\phi$-extended model).
Proposition 6.7 (Terminal knot-model invariants).
  1. 1.
  2. 2.
@@
By the compatibility condition (i) there exists $P_{1}$ $P_{2}$, $H$, and $\phi$ such that $M_{n}=P_{1}\otimes P_{2}$ and $G_{n}=H\otimes\phi^{-1}$. Therefore for Part 2 we can consider
@@
by Proposition A.1. Then, by compatibility condition (ii), we have
@@
Example 6.8 (Trivial terminal knot).
As in the standard case, the trivial knot does not change the Feynman–Kac model. At the other extreme is the adapted terminal knot, for which we discuss optimality in Section 6.4.
Example 6.9 (Adapted terminal knot).
6.3 Variance reduction and ordering
Theorem 6.10 (Variance difference with a terminal knot).
[…]
where $M_{n}=P_{1}\otimes P_{2}$ and $G_{n}=H\otimes\phi^{-1}$.
@@ 6.3 Variance reduction and ordering
from Proposition A.8 and noting $G_{n-1}^{\ast}=G_{n-1}$. Then since $M_{p}^{\ast}=M_{p}$ and $G_{p}^{\ast}=G_{p}$ for $p\in[0\,{:}\,n-1]$, we have $Q_{p}^{\ast}=Q_{p}$ for $p\in[1\,{:}\,n-1]$, and can state that $Q_{p,n}^{\ast}(G_{n}^{\ast}\cdot\bar{\varphi})=Q_{p,n}(G_{n}\cdot\bar{\varphi})$ for $p\in[0\,{:}\,n-1]$ and $Q_{n,n}^{\ast}=Q_{n,n}=\mathrm{Id}$ by definition. As such, combining with Proposition 6.7 (Part 1), we can state that $\hat{v}^{\ast}_{p,n}(\bar{\varphi})=\hat{v}_{p,n}(\bar{\varphi})$ for $p\in[0\,{:}\,n-1]$. Therefore, we can state that $\hat{\sigma}^{2}(\bar{\varphi})-\hat{\sigma}_{\ast}^{2}(\bar{\varphi})=\hat{v}_{n,n}(\bar{\varphi})-\hat{v}_{n,n}^{\ast}(\bar{\varphi})$.
@@
Corollary 6.11 (Variance reduction with a terminal knot).
  1. 1.
  2. 2.
[…]
@@
From Proposition 6.5 there exists Markov kernels $U,V$ such that $M_{n}^{\circ}=U\otimes V^{G_{n}\cdot\phi}$ and $G_{n}^{\circ}=V(G_{n}\cdot\phi)\otimes\phi^{-1}$. With $P_{1}=U$, $P_{2}=V^{G_{n}\cdot\phi}$, and $H=V(G_{n}\cdot\phi)$, we note that
@@
Equation 8 establishes that terminal knots applied to $\phi$-extended models with knots lead to an asymptotic variance ordering for test functions of the form $1\otimes\varphi$ when $\phi=|\varphi|$. Since $\mathcal{M}^{\circ}$ has this form, every knot in the assumed knot-model sequence reduces the variance of a test function of the form $1\otimes\varphi$. This follows from Proposition 4.1 (if a standard knot) or by (8) (if a terminal knot). Hence we can state that $\hat{\sigma}_{\circ}^{2}(1\otimes\varphi)\leq\hat{\sigma}_{\phi}^{2}(1\otimes\varphi)$ where $\hat{\sigma}_{\phi}^{2}$ is the asymptotic variance map for $\mathcal{M}^{\phi}$. Then from Proposition 6.2 we have $\hat{\sigma}_{\phi}^{2}(1\otimes\varphi)=\hat{\sigma}^{2}(\varphi)$ to complete the first part of the proof.
@@
  1. 1.
  2. 2.
Each knot in the sequences can be a terminal knots or a standard knot.
Theorem 6.13 (Variance ordering with terminal knots).
[…]
@@
6.4 Optimality of adapted terminal knots
Theorem 6.14 (Adapted terminal knot optimality).
@@ 6.4 Optimality of adapted terminal knots
First note that since $\mathcal{M}^{\circ}$ satisfies Part 2 of Definition 6.12 with respect to $\mathcal{M}$, by definition, $\mathcal{K}\ast\mathcal{M}^{\circ}$ also satisfies this condition. Then by Proposition 6.5 there exists Markov kernels $U,V$ such that $M_{n}^{\circ}=U\otimes V^{G_{n}\cdot\phi}$ and $G_{n}^{\circ}=V(G_{n}\cdot\phi)\otimes\phi^{-1}$. Let $\mathcal{K}=(n,R,K)$ and $\mathcal{R}=(n,\mathrm{Id},R)$, consider the model $\mathcal{M}^{\ast}=\mathcal{R}\ast\mathcal{K}\ast\mathcal{M}^{\circ}$, noting that $U=RK$ by compatibility. Hence $M_{n}^{\ast}=\mathrm{Id}\otimes R^{K(H)}K^{H}V^{G_{n}\cdot\phi}$ and $G_{n}^{\ast}=RK(H)\otimes\phi^{-1}$ where $H=V(G_{n}\cdot\phi)$ from the sequential application of the terminal knots. We can simplify the Markov kernel $M_{n}^{\ast}=\mathrm{Id}\otimes(RKV)^{G_{n}\cdot\phi}=\mathrm{Id}\otimes(UV)^{G_{n}\cdot\phi}$ using Proposition A.3 twice. The potential function simplifies to $G_{n}^{\ast}=UV(G_{n}\cdot\phi)\otimes\phi^{-1}$.
Now consider the model $\mathcal{M}^{\diamond}=\mathcal{K}^{\diamond}\ast\mathcal{M}^{\circ}$ where $\mathcal{K}^{\diamond}$ is the adapted kernel for $\mathcal{M}^{\circ}$. The adapted knot is $\mathcal{K}^{\diamond}=(n,\mathrm{Id},U)$ and hence $M_{n}^{\diamond}=\mathrm{Id}\otimes U^{H}V^{G_{n}\cdot\phi}=\mathrm{Id}\otimes(UV)^{G_{n}\cdot\phi}$ by Proposition A.3 and $G_{n}^{\ast}=U(H)\otimes\phi^{-1}=UV(G_{n}\cdot\phi)\otimes\phi^{-1}$. Hence, $M_{n}^{\diamond}=M_{n}^{\ast}$ and $G_{n}^{\diamond}=G_{n}^{\ast}$, so that we can conclude $\mathcal{M}^{\ast}=\mathcal{M}^{\diamond}$.
@@
Corollary 6.15 (Minimal variance from knotsets with terminal knot).
[…]
has asymptotic variance terms satisfying

### v2

@@ Appendix A Proof of main results
Proposition 6.5
@@
By compatibility condition (i), $\mathcal{M}^{\circ}=(M_{0:n}^{\circ},G_{0:n}^{\circ})$ satisfies $M_{n}^{\circ}=U_{1}\otimes V_{1}^{G_{n}\cdot\phi}$ and $G_{n}^{\circ}=V_{1}(G_{n}\cdot\phi)\otimes\phi^{-1}$, for some Markov kernels $U_{1}$ and $V_{1}$, reference model $\mathcal{M}=(M_{0:n},G_{0:n})$, and target function $\phi$. Let $\mathcal{M}^{\ast}=\mathcal{K}\ast\mathcal{M}^{\circ}=(M_{0:n}^{\ast},G_{0:n}^{\ast})$.
First case. If $\mathcal{K}$ is a non-terminal knot then $M_{n}^{\ast}=K^{G_{n-1}}U_{1}\otimes V_{1}^{G_{n}\cdot\phi}$ and $G_{n}^{\ast}=V_{1}(G_{n}\cdot\phi)\otimes\phi^{-1}$ for some Markov kernel $K$. Letting $U_{2}=K^{G_{n-1}}U_{1}$ and $V_{2}=V_{1}$ shows that $\mathcal{M}^{\ast}$ satisfies compatibility condition (i).
Second case. If $\mathcal{K}$ is a terminal knot then $M_{n}^{\ast}=R\otimes K^{V_{1}(G_{n}\cdot\phi)}V_{1}^{G_{n}\cdot\phi}$ and $G_{n}^{\ast}=KV_{1}(G_{n}\cdot\phi)\otimes\phi^{-1}$ for some Markov kernels $R$ and $K$ such that $RK=U_{1}$. By Proposition B.3, we can state $M_{n}^{\ast}=R\otimes(KV_{1})^{G_{n}\cdot\phi}$ and hence letting $U_{2}=R$ and $V_{2}=KV_{1}$ shows that $\mathcal{M}^{\ast}$ satisfies compatibility condition (i). ∎
Proposition 6.7
@@
By Definition 6.3 compatibility condition (i) there exists $P_{1}$ $P_{2}$, $H$, and $\phi$ such that $M_{n}=P_{1}\otimes P_{2}$ and $G_{n}=H\otimes\phi^{-1}$. Therefore for Part 2 we can consider
@@
by Proposition B.1. Then, by Definition 6.3 compatibility condition (ii), we have
@@
Theorem 6.12
@@
from Proposition B.8 and noting $G_{n-1}^{\ast}=G_{n-1}$. Then since $M_{p}^{\ast}=M_{p}$ and $G_{p}^{\ast}=G_{p}$ for $p\in[0\,{:}\,n-1]$, we have $Q_{p}^{\ast}=Q_{p}$ for $p\in[1\,{:}\,n-1]$, and can state that $Q_{p,n}^{\ast}(G_{n}^{\ast}\cdot\bar{\varphi})=Q_{p,n}(G_{n}\cdot\bar{\varphi})$ for $p\in[0\,{:}\,n-1]$ and $Q_{n,n}^{\ast}=Q_{n,n}=\mathrm{Id}$ by definition. As such, combining with Proposition 6.7 (Part 1), we can state that $\hat{v}^{\ast}_{p,n}(\bar{\varphi})=\hat{v}_{p,n}(\bar{\varphi})$ for $p\in[0\,{:}\,n-1]$. Therefore, we can state that $\hat{\sigma}^{2}(\bar{\varphi})-\hat{\sigma}_{\ast}^{2}(\bar{\varphi})=\hat{v}_{n,n}(\bar{\varphi})-\hat{v}_{n,n}^{\ast}(\bar{\varphi})$.
@@
Corollary 6.13
@@
From (repeated application of) Proposition 6.5 there exists Markov kernels $U,V$ such that $M_{n}^{\circ}=U\otimes V^{G_{n}\cdot\phi}$ and $G_{n}^{\circ}=V(G_{n}\cdot\phi)\otimes\phi^{-1}$. With $P_{1}=U$, $P_{2}=V^{G_{n}\cdot\phi}$, and $H=V(G_{n}\cdot\phi)$, we note that
@@
Equation (9) establishes that terminal knots applied to $\phi$-extended models with knots lead to an asymptotic variance ordering for test functions of the form $1\otimes\varphi$ when $\phi=|\varphi|$. Since $\mathcal{M}^{\circ}$ has this form, every knot in the assumed knot-model sequence reduces the variance of a test function of the form $1\otimes\varphi$. This follows from Proposition 4.1 (if a standard knot) or by (9) (if a terminal knot). Hence we can state that $\hat{\sigma}_{\circ}^{2}(1\otimes\varphi)\leq\hat{\sigma}_{\phi}^{2}(1\otimes\varphi)$ where $\hat{\sigma}_{\phi}^{2}$ is the asymptotic variance map for $\mathcal{M}^{\phi}$. Then from Proposition 6.2 we have $\hat{\sigma}_{\phi}^{2}(1\otimes\varphi)=\hat{\sigma}^{2}(\varphi)$ to complete the first part of the proof.
@@
Theorem 6.15
@@
Theorem 6.16
@@
First note that since $\mathcal{M}^{\circ}$ satisfies Part 2 of Definition 6.14 with respect to $\mathcal{M}$, by definition, $\mathcal{K}\ast\mathcal{M}^{\circ}$ also satisfies this condition. Then by (repeated application of) Proposition 6.5 there exists Markov kernels $U,V$ such that $M_{n}^{\circ}=U\otimes V^{G_{n}\cdot\phi}$ and $G_{n}^{\circ}=V(G_{n}\cdot\phi)\otimes\phi^{-1}$. Let $\mathcal{K}=(n,R,K)$ and $\mathcal{R}=(n,\mathrm{Id},R)$, consider the model $\mathcal{M}^{\ast}=\mathcal{R}\ast\mathcal{K}\ast\mathcal{M}^{\circ}$, noting that $U=RK$ by compatibility. Hence $M_{n}^{\ast}=\mathrm{Id}\otimes R^{K(H)}K^{H}V^{G_{n}\cdot\phi}$ and $G_{n}^{\ast}=RK(H)\otimes\phi^{-1}$ where $H=V(G_{n}\cdot\phi)$ from the sequential application of the terminal knots. We can simplify the Markov kernel $M_{n}^{\ast}=\mathrm{Id}\otimes(RKV)^{G_{n}\cdot\phi}=\mathrm{Id}\otimes(UV)^{G_{n}\cdot\phi}$ using Proposition B.3 twice. The potential function simplifies to $G_{n}^{\ast}=UV(G_{n}\cdot\phi)\otimes\phi^{-1}$.
Now consider the model $\mathcal{M}^{\diamond}=\mathcal{K}^{\diamond}\ast\mathcal{M}^{\circ}$ where $\mathcal{K}^{\diamond}$ is the adapted kernel for $\mathcal{M}^{\circ}$. The adapted knot is $\mathcal{K}^{\diamond}=(n,\mathrm{Id},U)$ and hence $M_{n}^{\diamond}=\mathrm{Id}\otimes U^{H}V^{G_{n}\cdot\phi}=\mathrm{Id}\otimes(UV)^{G_{n}\cdot\phi}$ by Proposition B.3 and $G_{n}^{\ast}=U(H)\otimes\phi^{-1}=UV(G_{n}\cdot\phi)\otimes\phi^{-1}$. Hence, $M_{n}^{\diamond}=M_{n}^{\ast}$ and $G_{n}^{\diamond}=G_{n}^{\ast}$, so that we can conclude $\mathcal{M}^{\ast}=\mathcal{M}^{\diamond}$.
@@
Corollary 6.17